OpenStudy (anonymous):

Solve. 12[ 10 - (5^2-6)^3/6]

OpenStudy (anonymous):

Following PEMDAS: Do Parenthesis first: \[5^2-10=15\]

OpenStudy (anonymous):

Oops; I should have subtracted 6 not 10: \[5^2-6=19\] followed by cubing: \[19^3=6859\]

OpenStudy (anonymous):

Division before subtraction: \[6859/6=1143.166...\]

OpenStudy (anonymous):

Now the subtraction in brackets (Parenthesis): \[10-1143.1666...=-1133.16666.\]

OpenStudy (anonymous):

and finally multiply by 12: \[12\times1133.1666...=-13598\]